OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -3x4 - 2x - 5 Neither Even Odd

OpenStudy (anonymous):

\[-3x^4-2x-5\]

OpenStudy (denonakavro):

It is neither according to graph

OpenStudy (irishboy123):

for even function f(-x) = f(x) for odd function f(-x) = -f(x)

OpenStudy (irishboy123):

apply the tests if it doesn't meet one of these tests, it is neither

OpenStudy (anonymous):

\[a_{n} = 3 \times(-3)^{n - 1}\]